A garden area is 30 ft long and 20 ft wide. A pat of uniform width is set around the edge. If the remaining garden is 400ft^2, what is the width of the path.

I'm not certain of my answer. It seems off, I am using the Quadratic Formula's for my answer. Am I on track??

Answer by stanbon(75887)

A garden area is 30 ft long and 20 ft wide. A path of uniform width is set inside the edge. If the remaining garden is 400ft^2, what is the width of the path.
==========
Probably.
Original Area = 30*20=600 sq ft
New Area = (30-2x)(20-2x)
EQUATION:
If width = x:
600-(30-2x)(20-2x)=400
Solve that quadratic for x.
========
Cheers,
Stan H.